# Taking the Guesswork Out of Language Model Evaluations

### v1

@@ 4 How Does Context Change Evaluation Conclusions?
Presence of context can improve agreement between evaluators.
We report agreement statistics (percentage agreement for majority label and Fleiss’ Kappa) for both autoraters and human evaluators in Tables 4 and 6 respectively. Significance testing is performed using a paired t-test and values significantly different than the NoCtxGen-NoCtxEval setting are indicated with a (p < 0.05). These statistics indicate that context-aware evaluation can significantly increase the agreement between evaluators, especially LM-based autoraters and sometimes human evaluators. This suggests that context is helpful in grounding evaluators together towards a more consistent judgement. To investigate this further, we compute human evaluator agreement when there was a difference of at least 1 in the total follow-up QAs satisfied by the two responses (reported in Table 8). These agreement rates are much higher, suggesting that the follow-up QAs can provide relatively more objective criteria to distinguish between responses.
Presence of context can change model rankings.
Based on the majority votes from autoraters and human evaluators, we report win rates for all model pairs in Tables 5 and 7 respectively. Note that we exclude those examples where there was no clear majority vote. The autorater win rates suggest that win rates can substantially vary across evaluation settings. Importantly, we find that relative rankings between pairs of models can flip in settings where evaluators are provided context versus when they are not. For instance, GPT-4o has a lower win rate than Gemini-1.5-Flash in the setting NoCtxGen-NoCtxEval but has a much higher win rate in the setting CtxGen-CtxEval. In general, these findings suggest that relative judgements between models can vary significantly when context is considered during evaluation.
@@
Presence of context can help evaluators use less surface-level criteria to judge responses.
Next, we investigate whether context changes the criteria used by evaluators to make judgements. To do this, we analyze the free-text justifications provided by human evaluators as well as autoraters. We automatically code these justifications into two categories: 1) surface-level criteria which includes criteria such as clarity, conciseness, style or formatting, tone and length, and 2) content-level criteria, which includes criteria such as relevance, correctness, completeness, level of detail and context adherence. We then automatically classify justifications into these two categories using GPT-4o. Trends across all evaluation settings for both human evaluators and autoraters are shown in Figure 5. We note that in context-aware evaluation settings, there is a lower percentage of justifications that are based on surface-level criteria and more that are based on content-level criteria.
@@ 5.1 Methodology
To investigate these implicit biases, we first define a list of contextual attributes and corresponding follow-up QAs (listed in Table 10 and based on Figure 2). For each attribute, we first filter those queries where this attribute is important to respond to the query and where the query is independent of the answer choices. For instance, a query such as "What is distillation in machine learning?" would be included for the attribute "User Expertise". Filtering is done using GPT-4o on 23,935 queries from all 5 datasets used in earlier experiments (prompt in Table 16). We then sample up to 1000 filtered queries randomly for each attribute and generate default responses (without context) for these queries using a candidate model.
@@ 5.2 Findings
First, we note that default responses are better catered to users who have a basic understanding of the topic of a query as opposed to experts, and the language they use lacks technical depth. Further, we find evidence of a WEIRD-like bias, where default responses are better aligned with western cultural contexts, middle-to-high income individuals, and young and middle-aged adults. Since context-agnostic evaluations might overlook disparities in how well default responses serve different contexts, we recommend future work to conduct similar analysis to discover implicit biases in default model responses.
@@ 6.1 Methodology
Models that are capable of adapting to various user contexts are useful for more individuals. Context-agnostic evaluations can miss out on capturing the adaptability of models to various contexts. We use our evaluation setting CtxGen-CtxEval to investigate how robustly models can adapt to different values of a contextual attribute. Similar to the analysis in §5, we use the contextual attributes and follow-up QAs defined in Table 10. However, we now generate adapted responses for every value of each contextual attribute where the follow-up QA (e.g., “What is your level of expertise on this topic?”) and a specific answer (e.g., “Complete beginner”) is provided to the generator.
@@ 7 Related Work
Language model-based autoraters hold great promise in improving the efficiency and reducing the cost of evaluations Chiang and Lee (2023). Prior work has identified that these autoraters are subject to biases such as preferring longer responses Dubois et al. (2024), preferring their own responses Panickssery et al. (2024) and others Zheng et al. (2023); Shen et al. (2023); Wang et al. (2024). Methods have been proposed to overcome some of these biases by providing clearer criteria during evaluation Liu et al. (2023); Saha et al. (2024). Our work proposes a simple recipe for improving the reliability of LM-based autoraters by providing context surrounding queries.
@@ 8 Limitations
The contextual attributes considered in our work only represent a sample of common information types that are missing in queries, which can be helpful to respond to a user’s query. The taxonomy presented in Figure 2 is not meant to be exhaustive and there are likely other types of context that are important.
@@ 10 Acknowledgments
First, we are thankful to all the human annotators who took time to participate in our studies. We also thank the following people for helpful discussions and comments: Sihao Chen, Doug Downey, Greg Durrett, Varsha Kishore, Nelson Liu, Nicholas Lourie, Vishakh Padmakumar, Valentina Pyatkin, Elizabeth Sieber, Luca Soldaini, David Wadden, Dan Weld, and Michael Zhang.
Appendix A Annotation Details
@@
We recruited a total of 1,085 participants through the Prolific crowdsourcing platform for the human evaluation studies. Annotators were required to have an approval rate of 99% and 500 prior successful submissions on the platform. They were also required to be fluent in English. Most participants came from the US and UK.
Annotation Setup.
For the main preference judgement task, we sampled a query and evaluation setting (from the three settings discussed in 3) randomly for each annotator. Annotators were also given a choice to skip a query for a different one, if they were unfamiliar with the topic of the original query. Each query-response pair was annotated by 3 different annotators and a single annotator was allowed to complete up to 3 examples. Before starting the task, annotators were required to complete a training task in the context-agnostic and context-aware evaluation settings, where they were given feedback about their ratings. Annotators were paid \$1.75 for each example at a rate of \$15 per hour, where they were allocated a total of 7 minutes per example.
@@
Screenshots of the annotation interfaces for context-agnostic and context-aware evaluation are provided in Figures 8 and 9 respectively.
@@ Appendix B Experimental Details
The prompt used for classifying queries into query types is provided in Table 11. The prompt for generating follow-up QAs for a query is provided in Table 12. For getting autorater judgements without and with context, the prompts are provided in Tables 13 and 14 respectively. Finally, the prompt to compute the number of constraints from the context met by a response is provided in Table 15. For the analyses presented in sections 5 and 6, we provide the prompt used to filter queries in Table 16, and the prompt used to rate responses on a Likert scale is in Table 17.
@@
Table 10: Contextual attributes and corresponding follow-up QA considered for the implicit context analysis presented in sections 5 and 6.
@@
Figure 8: Screenshot of the annotation interface used for context-agnostic evaluation.
Figure 9: Screenshot of the annotation interface used for context-aware evaluation.

### v2

@@ 5 How Does Context Change Evaluation Conclusions?
Presence of context improves agreement between evaluators.
We report percentage agreement rates for both autoraters and human evaluators in Table 4.333To calculate percentage agreement, we determine the majority label and then compute the percentage of judgments matching this label for each example. If all 3 human judgments choose the same option, this would be 100% agreement; if 2 of 3 humans choose the same option but one differs, this would be 66% agreement, and so on. Percentage agreement averages these scores across all examples. Significance testing is performed using a paired t-test and values significantly different than the NoCtxGen-NoCtxEval setting are indicated with a (p < 0.05). We find that context-aware evaluation can significantly increase the agreement between evaluators, especially LM-based autoraters and sometimes even human evaluators, suggesting that context is helpful in grounding evaluators together towards more consistent judgments.
While both autoraters and human evaluators generally show increased agreement when provided with context, the magnitude of improvement can vary. For example, autoraters show an improvement of roughly 3–10 points in agreement once context is introduced, whereas humans exhibit a more modest increase of about 3–5 points. This discrepancy suggests that LM-based evaluators may be more sensitive to context than human evaluators, potentially because they systematically focus on whether explicit constraints in the context are met. To investigate this further, we compute human agreement considering only those examples where the two responses differed by at least one in the number of follow-up QAs they satisfied (reported in Table 5). These agreement rates are much higher than the overall agreement rates in Table 4, suggesting that follow-up QAs can provide relatively more objective criteria to distinguish between responses.
Presence of context can change model win rates and flip benchmark rankings.
Based on the majority votes from autoraters and human evaluators, we report win rates for all model pairs in Table 6. Note that we exclude those examples where there was no clear majority vote. Overall, the autorater and human evaluator win rates suggest that win rates can substantially vary across evaluation settings. Importantly, we find that relative rankings between pairs of models can flip in settings where evaluators are provided context versus when they are not. For instance, in Table 6, GPT-4o has a lower win rate than Gemini-1.5-Flash in the setting NoCtxGen-NoCtxEval but has a much higher win rate in the setting CtxGen-CtxEval.
We observe similar trends between human evaluators and autoraters when examining magnitude of win rate shifts across evaluation settings, but note some key differences. For instance, while autoraters show a more significant flip favoring GPT-4o once context is provided, human evaluation shows a milder change, and is accompanied by an increased tie rate. When both responses incorporate the provided context sufficiently well, humans are less inclined to form a clear preference, resulting in a higher tie rate. On the other hand, autoraters rely on explicit checks for context fulfillment, which accentuates small differences in how each response incorporates context.
@@
Presence of context can help evaluators avoid reliance on surface-level criteria to judge responses.
Next, we investigate whether context changes the criteria used by evaluators to make judgments. To do this, we analyze the free-text justifications provided by human evaluators as well as autoraters. We automatically code these justifications into two categories: 1) surface-level criteria which includes criteria such as clarity, conciseness, style or formatting, tone and length, and 2) content-level criteria, which includes criteria such as relevance, correctness, completeness, level of detail and context adherence. We then automatically classify justifications into these two categories using GPT-4o. Trends across all evaluation settings for both human evaluators and autoraters are shown in Figure 4. We note that in context-aware evaluation settings, there is a lower percentage of justifications that are based on surface-level criteria and more that are based on content-level criteria.
@@ 6.1 Methodology
To investigate these implicit biases, we first define a list of contextual attributes and corresponding follow-up QAs (listed in Table 8 and based on Table 2). For each attribute, we first filter for those queries where this attribute is important to respond to the query and where the query is independent of the answer choices. For instance, a query such as "What is distillation in machine learning?" would correspond to the attribute "User Expertise". Filtering is done using GPT-4o on 23,935 queries from all 5 datasets used in earlier experiments (prompt in Table 14). We then sample up to 1000 filtered queries randomly for each attribute and generate default responses (without context) for these queries using a candidate model.
@@
For a subset of three contextual attributes ("Age Group", "Economic Context", "Cultural Context"), we also collect Likert ratings from human evaluators for the relevance of default responses to different possible values of each attribute (for e.g., possible values for "Age Group" -> [Children, Teenagers, Young adults, Middle-aged adults, Seniors]) for 75 queries (25 per contextual attribute). We then compute Spearman’s $p$ between human ratings for each query with the autorater judgements on the same query. Across the three contextual attributes, we obtain weak to moderate correlations (Age Group = 0.651, Cultural Context = 0.531, Economic Context = 0.276). We would like to emphasize that this task is highly subjective, which might affect the extent to which human ratings correlate with autorater ratings. At the same time, we find it encouraging that a subset of these attributes do have moderately positive correlations, which gives more credence to our findings below.
@@ 6.2 Findings
First, we note that default responses are better catered to users who have a basic understanding of the topic of a query as opposed to experts, and the language they use lacks technical depth. Further, our results suggest potential presence of a WEIRD-like bias, where default responses are better aligned with western cultural contexts, middle-to-high income individuals, and young and middle-aged adults. However, stronger correlations between human and model evaluators are required to strengthen this conclusion. Since context-agnostic evaluations might overlook disparities in how well default responses serve different contexts, we recommend future work to conduct similar analysis to discover implicit biases in default model responses.
@@ 7.1 Methodology
A common approach to adapting models to different user contexts is providing these contexts as part of a prompt; models that are capable of adapting to various user contexts are broadly more useful for more individuals. Context-agnostic evaluations can miss out on capturing the adaptability of models to various contexts. We use our evaluation setting CtxGen-CtxEval to investigate how robustly models can adapt to different values of a contextual attribute. Similar to the analysis in §6, we use the contextual attributes and follow-up QAs defined in Table 8. However, we now generate adapted responses for every value of each contextual attribute where the follow-up QA (e.g., “What is your level of expertise on this topic?”) and a specific answer (e.g., “Complete beginner”) is provided to the generator.
@@ 8 Related Work
Language model-based autoraters (also called LM-as-judge) hold great promise in improving the efficiency and reducing the cost of evaluations Chiang and Lee (2023). Prior work has identified that these autoraters are subject to biases such as preferring longer responses Dubois et al. (2024), preferring their own responses Panickssery et al. (2024) and others Zheng et al. (2023); Shen et al. (2023); Wang et al. (2024). Methods have been proposed to overcome some of these biases by providing clearer criteria during evaluation Liu et al. (2023); Saha et al. (2024). Our work proposes a simple recipe for improving the reliability of LM-based autoraters by providing context surrounding queries.
@@ 9 Limitations
The contextual attributes considered in our work only represent a sample of common information types that are missing in queries, which can be helpful to respond to a user’s query. The taxonomy presented in Table 2 is not meant to be exhaustive and there are likely other types of context that are important.
@@ 11 Acknowledgments
First, we thank all the human annotators who took time to participate in our studies. We also thank the following people for helpful discussions and comments: Sihao Chen, Doug Downey, Greg Durrett, Varsha Kishore, Nelson Liu, Nicholas Lourie, Vishakh Padmakumar, Valentina Pyatkin, Elizabeth Sieber, Luca Soldaini, David Wadden, Dan Weld, Michael Zhang and members of the Semantic Scholar research group at Ai2. Finally, we are grateful to the TACL reviewers and our action editor Omri Abend for their thoughtful comments and suggestions.
Appendix A Human Evaluation Details
@@
We recruited a total of 1,085 participants through the Prolific crowdsourcing platform for the human evaluation studies. Evaluators were required to have an approval rate of 99% and 500 prior successful submissions on the platform. They were also required to be fluent in English. Most participants came from the US and UK.
Evaluation Setup.
For the main preference judgment task, we sampled a query and evaluation setting (from the three settings discussed in 4) randomly for each evaluator. Evaluators were also given a choice to skip a query for a different one, if they were unfamiliar with the topic of the original query. Each query-response pair was judged by 3 different evaluators and a single evaluator was allowed to complete up to 3 examples. Before starting the task, evaluators were required to complete a training task in the context-agnostic and context-aware evaluation settings, where they were given feedback about their ratings. Evaluators were paid \$1.75 for each example at a rate of \$15 per hour, where they were allocated a total of 7 minutes per example.
@@
Screenshots of the evaluation interfaces for context-agnostic and context-aware evaluation are provided in Figures 7 and 8 respectively.
@@ Appendix B Experimental Details
The prompt used for classifying queries into query types is provided in Table 9. The prompt for generating follow-up QAs for a query is provided in Table 10. For getting autorater judgements without and with context, the prompts are provided in Tables 11 and 12 respectively. Finally, the prompt to compute the number of constraints from the context met by a response is provided in Table 13. For the analyses presented in sections 6 and 7, we provide the prompt used to filter queries in Table 14, and the prompt used to rate responses on a Likert scale is in Table 15.
@@
Table 8: Contextual attributes and corresponding follow-up QA considered for the implicit context analysis presented in Sections 6 and 7.
@@
Figure 7: Screenshot of the evaluation interface used for context-agnostic evaluation.
Figure 8: Screenshot of the evaluation interface used for context-aware evaluation.